# Position Paper: The Landscape and Challenges of HPC Research and LLMs

### 03 Feb 2024 v1

@@ 1 Introduction
Despite the recent success of LLMs in programming language-related tasks such as code generation and translation, there are no detailed studies of applying LLMs to HPC tasks. We conjecture that the unique characteristics of HPC pose distinct challenges in integrating LLM and HPC. To date, there has been no comprehensive exploration of the full potential or challenges of adapting and utilizing LLMs within the HPC domain. This represents a notable gap in our understanding, especially considering the rapid advancements in both fields. In this work, we aim to address this gap by systematically examining the reciprocal benefits that can arise from the intersection of LLMs and HPC. Further, we illustrate through a case study presented in Appendix B how tailored designs can facilitate mutual advantages between LLMs and HPC. This paper investigates ways in which LLMs can be tailored and optimized for HPC tasks while also addressing the challenges that accompany such integration.
@@ Appendix A HPC Downstream Tasks
Appendix B Case Studies and Potential Gaps
B.1 Instances of LLMs Enhancing HPC Performance
// Serial code for element-wise multiply
for (int i = 0; i < a.size(); i++) {
  a[i] = b[i] * c[i];
// Parallel version of the above serial code
#pragma omp parallel for
for (int i = 0; i < a.size(); i++) {
  a[i] = b[i] * c[i];
B.2 Instances of HPC Enhancing LLM Performance
HPC is pivotal in enhancing the performance of LLM across heterogeneous model deployments. The integration of HPC within environments characterized by diverse computational devices offers an ideal experimental platform for advancing Automated Machine Learning (AutoML) (Hutter et al., 2019) solutions tailored to LLMs, particularly within edge computing contexts. For instance, the efficient training or fine-tuning of dynamic weight-sharing models of different sizes (Muñoz et al., 2022), each selected to maximize the device’s capabilities where the model is deployed. Decentralized environments, such as Federated Learning (FL) (Li et al., 2019) frameworks, e.g., RaFL (Yu et al., 2022) and RaFFM (Yu et al., 2023) consider the resource-hungry characteristics of LLMs. For instance, RaFFM (Yu et al., 2023) employs salient parameter prioritization to create the abstraction of an extensive search space of efficient weight-sharing models with fewer parameters that can be efficiently extracted and deployed to an FL client with enough resources to optimize the received model while maximizing its capabilities. Low latency and high throughput are crucial for the effectiveness of this advanced FL framework.
Sophisticated solutions must efficiently serve the optimized models in environments with diverse computing platforms. Finally, HPC challenges are also present during the LLM inference and generation stages, in which the model, either deployed at the edge or using remote inference, must satisfy the user’s efficiency and accuracy requirements. To respond to some of these challenges, the open-source community has proposed vLLM (Kwon et al., 2023b). This library integrates with other popular libraries, e.g., Huggingface (Wolf et al., 2020), to improve the efficiency of existing LLMs. For instance, vLLM implements the PagedAttention algorithm (Kwon et al., 2023a), improving the management of the Key-Value cache in Transformer-based architectures, resulting in more efficient models with higher throughput. HPC can play a significant role in these solutions.

### 07 Feb 2024 v3

# The Landscape and Challenges of HPC Research and LLMs

@@ 1 Introduction
Despite the recent success of LLMs in programming language-related tasks such as code generation and translation, there are no detailed studies of applying LLMs to HPC tasks. We conjecture that the unique characteristics of HPC pose distinct challenges in integrating LLM and HPC. To date, there has been no comprehensive exploration of the full potential or challenges of adapting and utilizing LLMs within the HPC domain. This represents a notable gap in our understanding, especially considering the rapid advancements in both fields. In this work, we aim to address this gap by systematically examining the reciprocal benefits that can arise from the intersection of LLMs and HPC. Further, we illustrate through a case study presented in Section 6 how tailored designs can facilitate mutual advantages between LLMs and HPC. This paper investigates ways in which LLMs can be tailored and optimized for HPC tasks while also addressing the challenges that accompany such integration.
@@ 5 Ethical Considerations
6 Case Studies and Potential Gaps
6.1 Instances of LLMs Enhancing HPC Performance
// Serial code for element-wise multiply
for (int i = 0; i < a.size(); i++) {
  a[i] = b[i] * c[i];
// Parallel version of the above serial code
#pragma omp parallel for
for (int i = 0; i < a.size(); i++) {
  a[i] = b[i] * c[i];
6.2 Instances of HPC Enhancing LLM Performance
HPC is crucial for efficiently enhancing LLM performance. For instance, in Federated Learning (FL) (Li et al., 2019) setups, HPC environments with heterogeneous devices can benefit from the efficient training or fine-tuning of dynamic weight-sharing models of different sizes (Muñoz et al., 2022), each selected to maximize the device’s capabilities where the model is deployed. RaFFM (Yu et al., 2023) is an example of an FL framework that considers the resource-hungry characteristics of LLMs (Yu et al., 2023). This FL framework employs salient parameter prioritization to create the abstraction of an extensive search space of efficient weight-sharing models with fewer parameters that can be efficiently extracted and deployed to an FL client with enough resources to optimize the received model while maximizing its capabilities. Low latency and high throughput are crucial for the effectiveness of this advanced FL framework.
Sophisticated solutions must efficiently serve the optimized models in environments with diverse computing platforms. Finally, HPC challenges are also present during the LLM inference and generation stages, in which the model, either deployed at the Edge or using remote inference, must satisfy the user’s efficiency and accuracy requirements. To respond to some of these challenges, the open-source community has proposed vLLM (Kwon et al., 2023b). This library integrates with other popular libraries, e.g., Huggingface (Wolf et al., 2020), to improve the efficiency of existing LLMs. For instance, vLLM implements the PagedAttention (Kwon et al., 2023a) algorithm, improving the management of the Key-Value cache in Transformer-based architectures, resulting in more efficient models with higher throughput. HPC can play a significant role in these solutions.